if x=5y-1 and 2x+5y=-32 what is the value for y
a.-2
b.2
c.1
d.-1

Answer by tutorcecilia(2144)

x=5y-1 [substitute "5y-1" for x in the second equation]
2(5y-1)+5y=-32 [simplify]
10y-2+5y=-32
15y-2=-32 [solve for y]
15y=-32+2
15y=-30
y=-2
.
check by plugging y into either equation and solve for x
x=5y-1
x=5(-2)-1
x=-10-1
x=-11
.
-11=5(-2)-1
-11=-10-1
-11=-11 [checks out]